A car approaches a stationary police car at 36 m/s . The frequency of the siren (relative to the police car ) is 500 Hz. What is the frequency (in Hz) heard by an observer in the moving car as he approaches the police car? (Assume the velocity of sound in air is 343 m/ s.)


1
Expert's answer
2021-06-29T10:00:43-0400

The Doppler Effect equation says

"f=f_0\frac{c+v}{c}\\
=500\:\rm Hz\frac{(343+36) \:m/s}{343\: m/s}=552\: Hz"